2x - 7√x - 30 = 0.
Solve by factoring.

Answer by stanbon(75887)

2x - 7√x - 30 = 0.
(2sqrtx + 5) (sqrt x - 6)=0
sqrtx = -5/2 or sqrtx =6
x=(25/4) or x=36
Checking these solutions:
2(25/4)-7(5/4)-30=0
50/4-35/4-120/4= 0
Not true so 25/4 is an extraneous solution.
2(36)-7(6)-30=0
0=0
So, x=36 is the only solution.
Cheers,
Stan H.